OpenStudy (anonymous):

Determine if triangle EST with coordinates R(3,4) S(5,5) and T (6,1) is a right triangle. Use evidence to support your claim. If it is not a right triangle what changes can be made to make it a right triangle? Be specific

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

1) use the distance formula to find each of the sides. 2) do these sides fit into a²+b²=c² ?

OpenStudy (anonymous):

Can u please help me do this I'm so confused

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf d= \sqrt{(y_1-y_2)^{2}+(x_1-x_2)^{2}} }\) that is the distance formula.

OpenStudy (solomonzelman):

Can you find each side using this formula ?

OpenStudy (anonymous):

I don't know how to use the formula @SolomonZelman sorry

OpenStudy (solomonzelman):

\(\LARGE\color{blue}{ \bf d= \sqrt{(y_1-y_2)^{2}+(x_1-x_2)^{2}} }\) R(3,4) S(5,5) RS= \(\Large\color{blue}{ \bf RS= \sqrt{(4-5)^{2}+(3-5)^{2}}=\sqrt{(1)^{2}+(2)^{2}} }\) \(\Large\color{blue}{ \bf RS= \sqrt{(1)^{2}+(2)^{2}}=\sqrt{1+2} =\sqrt{3} }\) I found RS, can you find the remaining sides similarly ?

OpenStudy (solomonzelman):

My bad \[\sqrt{(1)^2+(-2)^2} =\sqrt{1+4} =\sqrt{5} \]

OpenStudy (anonymous):

5-6)+5-1 not sure

OpenStudy (solomonzelman):

(5,5) and T (6,1) \[ST=\sqrt{ (5-1)^2+(5-6)^2}\]

OpenStudy (solomonzelman):

See how I am plugging the points in ?

OpenStudy (anonymous):

Ya but still sort of confuse

OpenStudy (solomonzelman):

Can you simplify the square root of what I wrote ?

OpenStudy (anonymous):

I don't know if this is right but I got 2.34

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

don't simplify the root, tell me what you get in terms of square root, and simplify the result as much as possible.

OpenStudy (anonymous):

the 6 can become a 3 @SolomonZelman

OpenStudy (solomonzelman):

I don;t get what you are saying right now

OpenStudy (anonymous):

From St= (5-1)^2+(5-6)^2 the 6 could be simplified to a 3

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

Please help so confused